# On the Expressive Power of Modern Hopfield Networks

### v1

@@
On the Expressive Power of Modern Hopfield Networks
Xiaoyu Li xli216@stevens.edu. Stevens Institute of Technology.    Yuanpeng Li physicsyp@gmail.com. Beijing Normal University.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at UC Berkeley.
Modern Hopfield networks (MHNs) have emerged as powerful tools in deep learning, capable of replacing components such as pooling layers, LSTMs, and attention mechanisms. Recent advancements have enhanced their storage capacity, retrieval speed, and error rates. However, the fundamental limits of their computational expressiveness remain unexplored. Understanding the expressive power of MHNs is crucial for optimizing their integration into deep learning architectures. In this work, we establish rigorous theoretical bounds on the computational capabilities of MHNs using circuit complexity theory. Our key contribution is that we show that MHNs are $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$. Hence, unless $\mathsf{TC}^{0}=\mathsf{NC}^{1}$, a $\mathrm{poly}(n)$-precision modern Hopfield networks with a constant number of layers and $O(n)$ hidden dimension cannot solve $\mathsf{NC}^{1}$-hard problems such as the undirected graph connectivity problem and the tree isomorphism problem. We also extended our results to Kernelized Hopfield Networks. These results demonstrate the limitation in the expressive power of the modern Hopfield networks. Moreover, Our theoretical analysis provides insights to guide the development of new Hopfield-based architectures.
@@ 1 Introduction
Hopfield networks [25], initially introduced as associative memories capable of storing and retrieving patterns, have undergone significant advancements in recent years. These developments led to the emergence of modern Hopfield networks [59]. Modern Hopfield networks (MHNs) address the limitations of their predecessors, particularly in terms of storage capacity, retrieval speed, and error rates. An important feature of MHNs is their ability to function as specialized components within deep networks, integrating memory capabilities and supporting diverse functionalities. For instance, MHNs can replace conventional layers such as pooling layers, permutation-equivariant layers [19, 60], GRU [13], LSTM [24, 27], and attention mechanisms [69, 5].
@@
In this work, we present a comprehensive theoretical investigation into the circuit complexity bounds of MHNs. Our approach involves analyzing the architecture of MHNs and the computational complexity of its components, such as the Hopfield layer. Hence, we show that uniform $\mathsf{TC}^{0}$ circuits can efficiently simulate these models.
Our main contributions can be outlined as follows:
  • We prove that any $\operatorname{poly}(n)$-precision modern Hopfield networks or kernelized Hopfield networks with constant-depth and $O(n)$ hidden dimension is in $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ circuit family (Theorem 4.6 and 5.4).
  • We prove that unless $\mathsf{TC}^{0}=\mathsf{NC}^{1}$, a $\operatorname{poly}(n)$-precision modern Hopfield networks or kernelized Hopfield networks with constant layers, $O(n)$ hidden dimension cannot solve $\mathsf{NC}^{1}$-hard problems such as the undirected graph connectivity problems and tree isomorphism problems (Theorems 6.1, 6.2, 6.3 and 6.4).
@@
In Section 2, we provide an overview of the related works. In Section 3, we introduce the notations and definitions needed for modern Hopfield networks. In Section 4, we analyze the circuit complexity for modern Hopfield networks. In Section 5, we focus on the circuit complexity results of kernelized Hopfield networks. In Section 6, we discuss the hardness results of the modern Hopfield networks. Finally, in Section 7, we summarize our theoretical results in the conclusion. We defer more preliminaries and technique details in Appendix.
@@ 2 Related Works
In this section, we introduce the related works of the Hopfield models and the complexity of Transformers. In Appendix A, we introduce more related works.
2.1 Theory of Modern Hopfield Networks
Modern Hopfield Models have demonstrated both empirical success and theoretical value, providing a framework for understanding transformer attention mechanisms and Transformer architectures. [31, 70] propose a unified framework for analyzing and deriving modern Hopfield models using entropic regularizers. This framework includes a range of sparse variants, such as sparse and generalized sparse models, and also integrates the standard modern Hopfield model [59] as a special case. Despite these advancements, the modern Hopfield paradigm remains incomplete, lacking efficient implementations and variants, as highlighted by [31] To address these limitations, [21] propose a principled nonparametric approach for developing efficient variants of modern Hopfield models, including linear, top-K, and random feature-based versions. This study focuses on advancing research to create more efficient models. This study aims to push the boundaries of research in developing more efficient models, emphasizing their potential to shape the future of Hopfield-driven designs.
2.2 Complexity and Neural Network
@@ 3 Preliminaries
In Section 3.1, we define useful notations for defining modern Hopfield networks. In Section 3.2, we introduce some components of modern Hopfield networks. In Section 3.3, we state the basic definitions and architectures of kernelized Hopfiled networks. The background knowledge about circuit complexity and float point number is introduced in Appendix B.
@@ 3.1 Notations
We define :=$\mathbb{N}:=\{0,1,2,\ldots\}$ as the set of natural numbers and $\mathbb{R}$ as the set of real numbers. For any positive integer $n$, $[n]$ represents the set $\{1,2,\ldots,n\}$. The vector $\mathbf{1}_{n}$ denotes an $n$-dimensional vector where all entries are ones. Given a matrix $A\in\mathbb{R}^{m\times n}$, $A_{i,j}$ refers to the element in the $i$-th row and $j$-th column, while $A^{\top}$ represents the transpose of $A$. The set $\{0,1\}^{*}$ represents all binary strings of finite length. Specifically, for $x_{i}\in\{0,1\}^{*}$, $x_{i}$ represents a finite binary string, with each bit being either 00 or $1$. A language $L$ is characterized as a subset of $\{0,1\}^{*}$.
@@ 3.2 Modern Hopfield Network
In [59], the Modern Hopfield Model is introduced with a specific energy function $E$ and retrieval dynamics $\mathcal{T}$, which are further incorporated into deep learning frameworks due to their relation to transformer attention mechanisms [69]. This approach enhances performance and provides a theoretical guarantee of exponential memory capacity. The energy function is defined as:
$E(x)=-\mathrm{lse}(\beta,\Xi^{\top}x)+\frac{1}{2}\langle x,x\rangle,$
where the retrieval dynamics is given by
@@
The floating-point number $\mathbb{F}_{p}$ and their operations in our computational framework are formally defined in Appendix B.2. The Hopfield attention marix is used to compute a single Hopfield layer.
@@
$\mathsf{MHN}(R)=f_{m}\circ\mathsf{Hop}_{m}(f_{m-1}\circ\mathsf{Hop}_{m-1}(\cdots f_{1}\circ\mathsf{Hop}_{1}(f_{0}(R),Y_{1})\cdots,Y_{m-1}),Y_{m})\in\mathbb{F}_{p}^{n\times d},$
@@ 3.3 Kernelized Hopfield Networks
$\mathcal{T}_{\Phi}(x):=\Xi\cdot\mathrm{Softmax}(\beta\mathcal{K}(\Xi,x)),\quad E_{\mathcal{K}}(x)=\frac{1}{2}\mathcal{K}(x,x)+\mathrm{lse}(\beta,\mathcal{K}(\Xi,x)),$
@@
Let $R\in\mathbb{F}_{p}^{n\times d}$ be the set of query (state) patterns, and $Y\in\mathbb{F}_{p}^{n\times d}$ be the set of key (stored) patterns. Let $W_{Q}\in\mathbb{F}_{p}^{d\times D_{\Phi}},W_{K}\in\mathbb{F}_{p}^{d\times D_{\Phi}}$, and $W_{V}\in\mathbb{F}_{p}^{d\times d}$ be learnable projection matrices. Consider a feature map $\Phi:\mathbb{F}_{p}^{D_{\Phi}}\rightarrow\mathbb{F}_{p}^{D_{\Phi}}$ associated with a kernel function $\mathcal{K}:\mathbb{F}_{p}^{D_{\Phi}}\times\mathbb{F}_{p}^{D_{\Phi}}\rightarrow\mathbb{F}_{p}$ defined by $\mathcal{K}(\cdot,\cdot):=\langle\Phi(\cdot),\Phi(\cdot)\rangle$. Let $\beta>0$ be a scaling parameter. According to Definition 3.1, the kernelized Hopfield attention matrix $A\in\mathbb{F}_{p}^{n\times n}$ is defined by, for every $i,j\in[n]$,
$A_{i,j}:=\exp(\beta\cdot\langle\Phi(R_{i}W_{Q}),\Phi(Y_{j}W_{K})\rangle)$
@@
Note that the softmax is applied row-wise, $\langle\Phi(RW_{Q}),\Phi(YW_{K})\rangle_{i,j}=\langle\Phi(R_{i}W_{Q}),\Phi(Y_{j}W_{K})\rangle$. We set $D:=\operatorname{diag}(\beta\cdot A1_{n})\in\mathbb{F}_{p}^{n\times n}$. Then, based on Definition 3.1, we define the i-th kernelized Hopfield layer as
$\mathsf{KHop}_{i}(R,Y_{i}):=D^{-1}AY_{i}W_{V}$
@@
$\mathsf{KHN}(R)=f_{m}\circ\mathsf{KHop}_{m}(f_{m-1}\circ\mathsf{KHop}_{m-1}(\cdots f_{1}\circ\mathsf{KHop}_{1}(f_{0}(X),Y_{1})\cdots,Y_{m-1}),Y_{m})\in\mathbb{F}_{p}^{n\times d},$
@@ 4 Complexity of Modern Hopfield Networks
This section explores key results regarding the circuit complexity of the computations involved in modern Hopfield networks. We begin with an analysis of the Hopfield attention matrix in Section 4.1. In Section 4.2, we delve into the computation of a single Hopfield layer. Section 4.3 extends the discussion to other components beyond the Hopfield layer. In Section 4.4, we examine the modern Hopfield network. Finally, Section 4.5 presents our main result: the circuit complexity bounds for modern Hopfield networks. These findings form the basis for our main theorem on the expressive power of Hopfield networks. In Appendix B.1, we provide fundamental definitions from circuit complexity theory.

### v2

# Modern Hopfield Networks Require Chain-of-Thought to Solve $\mathsf{NC}^{1}$-Hard Problems

@@
Modern Hopfield Networks Require Chain-of-Thought to Solve $\mathsf{NC}^{1}$-Hard Problems
Yang Cao and Xiaoyu Li and Yuanpeng Li and Yingyu Liang and Zhenmei Shi and Zhao Song Wyoming Seminary.University of New South Wales.Independent Researcher.The University of Hong Kong.University of Wisconsin-Madison.magic.linuxkde@gmail.com. Simons Institute for the Theory of Computing, University of California, Berkeley.
Modern Hopfield Networks (MHNs) have emerged as powerful components in deep learning, serving as effective replacements for pooling layers, LSTMs, and attention mechanisms. While recent advancements have significantly improved their storage capacity and retrieval efficiency, their fundamental theoretical boundaries remain underexplored. In this paper, we rigorously characterize the expressive power of MHNs through the lens of circuit complexity theory. We prove that $\mathrm{poly}(n)$-precision MHNs with constant depth and linear hidden dimension fall within the $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ complexity class. Consequently, assuming $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$, we demonstrate that these architectures are incapable of solving $\mathsf{NC}^{1}$-hard problems, such as undirected graph connectivity and tree isomorphism. We further extend these impossibility results to Kernelized Hopfield Networks. However, we show that these limitations are not absolute: we prove that equipping MHNs with a Chain-of-Thought (CoT) mechanism enables them to transcend the $\mathsf{TC}^{0}$ barrier, allowing them to solve inherently serial problems like the word problem for the permutation group $S_{5}$. Collectively, our results delineate a fine-grained boundary between the capabilities of standard MHNs and those augmented with reasoning steps.
@@ 1 Introduction
Hopfield networks [20], initially introduced as associative memories capable of storing and retrieving patterns, have undergone significant advancements in recent years. These developments led to the emergence of Modern Hopfield Networks [40]. Modern Hopfield Networks (MHNs) address the limitations of their predecessors, particularly in terms of storage capacity, retrieval speed, and error rates. An important feature of MHNs is their ability to function as specialized components within deep networks, integrating memory capabilities and supporting diverse functionalities. For instance, MHNs can replace conventional layers such as pooling layers, permutation-equivariant layers [14, 41], GRU [10], LSTM [19, 22], and attention mechanisms [46, 4]. Their energy-based formulation allows for rapid and efficient pattern retrieval, making them a compelling alternative to traditional memory structures in deep learning.
@@
In this work, we present a comprehensive theoretical investigation into the circuit complexity bounds of MHNs. Our approach involves a detailed analysis of the MHN architecture and the computational complexity of its constituent components, such as the Hopfield layer. Consequently, we demonstrate that these models can be efficiently simulated by uniform $\mathsf{TC}^{0}$ circuits.
Our main contributions can be summarized as follows:
  • We establish that any $\mathrm{poly}(n)$-precision Modern Hopfield Network or Kernelized Hopfield Network with constant depth and $O(n)$ hidden dimension resides within the $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ complexity class (Theorems 4.8 and 5.4).
  • We demonstrate that, assuming $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$, $\mathrm{poly}(n)$-precision Modern or Kernelized Hopfield Networks with constant layers and $O(n)$ hidden dimension are fundamentally incapable of solving $\mathsf{NC}^{1}$-hard problems. We specifically identify undirected graph connectivity and tree isomorphism as problems beyond the reach of these architectures (Theorems 6.9, 6.10, 6.11 and 6.12).
  • We prove that these computational limitations are not absolute and can be overcome by the Chain-of-Thought (CoT) mechanism. Specifically, we show that a Modern Hopfield Network with $n$-step CoT can solve the word problem for the permutation group $S_{5}$ (an $\mathsf{NC}^{1}$-complete problem), a task that is provably impossible for standard constant-depth MHNs (Theorem 7.4).
@@
In Section 2, we provide an overview of the related works. In Section 3, we introduce the notations and definitions needed for Modern Hopfield Networks. In Section 4, we analyze the circuit complexity for Modern Hopfield Networks. In Section 5, we focus on the circuit complexity results of Kernelized Hopfield Networks. In Section 6, we discuss the hardness results of the Modern Hopfield Networks. In Section 7, we provide our the results for Modern Hopfield Networks with Chain-of-Thought mechanism. Finally, in Section 8, we summarize our theoretical results in the conclusion.
@@ 2 Related Works
In this section, we introduce the related works of the Hopfield models and the complexity of Transformers.
Theory of Modern Hopfield Networks
The Hopfield network, introduced by John Hopfield [20, 21, 28, 29, 30], was originally conceived as a model for content-addressable memory. This network operates by encoding patterns through the connections between neurons, enabling associative recall. By fixing specific inputs and allowing the network to evolve dynamically, it minimizes an energy function to retrieve stored patterns. A notable characteristic of Hopfield networks is the overlapping nature of different patterns, where many patterns share common neurons. Recent advancements in Modern Hopfield Networks have showcased both theoretical significance and practical applications. These models have proven instrumental in offering insights into transformer attention mechanisms and the architecture of Transformers. For instance, [24, 47] introduced a comprehensive framework for analyzing and deriving modern Hopfield models, employing entropic regularizers to unify diverse variants. This framework encompasses sparse and generalized sparse models and includes the standard modern Hopfield network [40] as a particular instance. Moreover, contemporary research efforts [24, 48, 18, 15, 16, 23] continue to explore innovative directions for improving the efficiency and scalability of Hopfield-based models. These studies underline the transformative potential of Hopfield-driven methodologies in shaping future designs, highlighting their adaptability to complex computational tasks and their integration into broader machine learning paradigms.
Modern Hopfield Networks and Applications in Deep Learning.
Classical Hopfield networks [20, 21, 28, 29, 30] have long been recognized for their ability to emulate human brain associative memory, primarily focusing on the storage and retrieval of memory patterns. Originally limited by linear memory capacity [28], these networks now achieve exponential storage capabilities [11] and kernelized [47], allowing for efficient handling of vast memory patterns. Architectural innovations [17, 43, 12] have integrated Hopfield networks into models like Transformers, serving as advanced attention mechanisms [46]. Their biological plausibility [31, 30] has been a key focus, aligning computational models more closely with human associative memory. These developments have expanded applications across various fields, including tabular data learning [50], drug discovery [44], immunology [49], time series forecasting [47], reinforcement learning [39], and large foundation models [12], demonstrating their versatility in addressing complex problems.
Complexity and Neural Network
@@
Limitations of Transformer Architectures.
While Transformers excel in natural language processing, their capabilities in mathematical computation remain constrained [5]. To understand these limitations, researchers have examined two types of Transformers: (1) average-head attention, which assigns 1 to the largest probability entry and 0 to others, and (2) softmax-attention, which computes probabilities as $\mathsf{Softmax}(X)=\operatorname{diag}(\exp(X)\cdot{\bf 1}_{n})^{-1}\cdot\exp(X)$. Average-head Transformers can recognize languages beyond $\mathsf{AC}^{0}$, but they are confined to $\mathsf{TC}^{0}$, as shown by [38]. Similarly, softmax-attention Transformers are also within $\mathsf{TC}^{0}$ [32]. Extending this, [37] formalize a generalized similarity function $s$ and demonstrate that softmax-attention Transformers fall under $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. Through a first-order logic framework with $\mathsf{MAJORITY}$ quantifiers ($\mathsf{FOM}$[25], [36] show that these Transformers can be simulated in $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$. [6] further enhances precision, proving that Transformers with minimal error ($2^{-O(\mathrm{poly}(n))}$) also belong to this class. For practical problems, [13] reveal that unless $\mathsf{TC}^{0}=\mathsf{NC}^{1}$, log-precision Transformers cannot solve arithmetic, equation-solving, or Context-Free Grammar (CFG) membership tasks [42]. These findings explain why Transformers often struggle with mathematical reasoning. Recently, [7] studied the circuit complexity bounds of RoPE-based transformer architectures.
@@ 3 Preliminaries
In Section 3.1, we define useful notations for defining Modern Hopfield Networks. In Section 3.2, we introduce the definition of floating-point numbers. In Section 3.3, we introduce some components of Modern Hopfield Networks. In Section 3.4, we state the basic definitions and architectures of kernelized Hopfiled networks. The background knowledge about circuit complexity is introduced in Section 4.1.
@@ 3.1 Notations
We define $\mathbb{N}:=\{0,1,2,\ldots\}$ as the set of natural numbers and $\mathbb{R}$ as the set of real numbers. For any positive integer $n$, $[n]$ represents the set $\{1,2,\ldots,n\}$. The vector $\mathbf{1}_{n}$ denotes an $n$-dimensional vector where all entries are ones. Given a matrix $A\in\mathbb{R}^{m\times n}$, $A_{i,j}$ refers to the element in the $i$-th row and $j$-th column, while $A^{\top}$ represents the transpose of $A$. The set $\{0,1\}^{*}$ represents all binary strings of finite length. Specifically, for $x_{i}\in\{0,1\}^{*}$, $x_{i}$ represents a finite binary string, with each bit being either 0 or $1$. A language $L$ is characterized as a subset of $\{0,1\}^{*}$.
3.2 Floating-Point Numbers
Definition 3.1 (Floating-point number, Definition 9 in [6]).
Definition 3.2 (Rounding, Definition 9 in [6]).
Lemma 3.3 (Basic floating-point operations in $\mathsf{TC}^{0}$, Lemmas 10, 11 in [6]).
Part 1. Let $x_{1},x_{2}$ be two $p$-bits floating point numbers. The arithmetic operations of addition, multiplication, division, and comparison of $x_{1}$ and $x_{2}$, as described in [6], can be executed using a $\mathsf{DLOGTIME}$-uniform threshold circuit characterized by constant depth, and a size bounded by $\operatorname{poly}(n)$. The maximum depth of the circuit required is denoted by $d_{\mathrm{std}}$.
Part 2. Let $x_{1},\ldots,x_{n}$ be $n$ p-bit floating-point number. Iterative multiplication of $x_{1},\ldots,x_{n}$ can be executed with a $\mathsf{DLOGTIME}$-uniform threshold circuit characterized by constant depth and a size bounded by $\operatorname{poly}(n)$. The circuit depth for this operation is indicated as $d_{\otimes}$.
Part 3. Let $x_{1},\ldots,x_{n}$ be $n$ p-bit floating-point number. Iterative addition of $x_{1},\ldots,x_{n}$, where rounding is applied after the summation is completed, can be achieved using a $\mathsf{DLOGTIME}$-uniform threshold circuit characterized by constant depth, and a size bounded by $\operatorname{poly}(n)$. The depth required for this operation is represented by $d_{\oplus}$.
This lemma shows that basic arithmetic operations can be efficiently executed in $\mathsf{TC}^{0}$ when operating on floating-point numbers with precision $p\leq\operatorname{poly}(n)$.
Lemma 3.4 (Computing $\exp$ in $\mathsf{TC}^{0}$, Lemma 12 in [6]).
Lemma 3.5 (Approximating square root in $\mathsf{TC}^{0}$, Lemma 12 in [6]).
@@ 3.3 Modern Hopfield Network
In [40], the Modern Hopfield Model is introduced with a specific energy function $E$ and retrieval dynamics $\mathcal{T}$, which are further incorporated into deep learning frameworks due to their relation to transformer attention mechanisms [46]. This approach enhances performance and provides a theoretical guarantee of exponential memory capacity. The energy function is defined as: $E(x)=-\mathrm{lse}(\beta,\Xi^{\top}x)+\frac{1}{2}\langle x,x\rangle,$ where the retrieval dynamics is given by
@@
$\mathbb{F}_{p}$ is the set of all $p$-bit floating-point numbers. The floating-point number set $\mathbb{F}_{p}$, and their operations in our computational framework are formally defined in Section 3.2. The Hopfield attention matrix is used to compute a single Hopfield layer.
@@
$\mathsf{MHN}(R)=$ $~f_{m}\circ\mathsf{Hop}_{m}(f_{m-1}\circ\mathsf{Hop}_{m-1}(\cdots f_{1}\circ\mathsf{Hop}_{1}(f_{0}(R),Y_{1})\cdots,Y_{m-1}),Y_{m}),$
@@ 3.4 Kernelized Hopfield Networks
$\mathcal{T}_{\Phi}(x):=$ $~\Xi\cdot\mathrm{Softmax}(\beta\mathcal{K}(\Xi,x))$
$\quad E_{\mathcal{K}}(x):=$ $~\frac{1}{2}\mathcal{K}(x,x)+\mathrm{lse}(\beta,\mathcal{K}(\Xi,x)),$
@@
Let $R\in\mathbb{F}_{p}^{n\times d}$ be the set of query (state) patterns, and $Y\in\mathbb{F}_{p}^{n\times d}$ be the set of key (stored) patterns. Let $W_{Q}\in\mathbb{F}_{p}^{d\times D_{\Phi}},W_{K}\in\mathbb{F}_{p}^{d\times D_{\Phi}}$, and $W_{V}\in\mathbb{F}_{p}^{d\times d}$ be learnable projection matrices. Consider a feature map $\Phi:\mathbb{F}_{p}^{D_{\Phi}}\rightarrow\mathbb{F}_{p}^{D_{\Phi}}$ associated with a kernel function $\mathcal{K}:\mathbb{F}_{p}^{D_{\Phi}}\times\mathbb{F}_{p}^{D_{\Phi}}\rightarrow\mathbb{F}_{p}$ defined by $\mathcal{K}(\cdot,\cdot):=\langle\Phi(\cdot),\Phi(\cdot)\rangle$. Let $\beta>0$ be a scaling parameter. According to Definition 3.6, the kernelized Hopfield attention matrix $A\in\mathbb{F}_{p}^{n\times n}$ is defined by, for every $i,j\in[n]$, $A_{i,j}:=\exp(\beta\cdot\langle\Phi(R_{i}W_{Q}),\Phi(Y_{j}W_{K})\rangle)$
@@
Note that the softmax is applied row-wise, $\langle\Phi(RW_{Q}),\Phi(YW_{K})\rangle_{i,j}=\langle\Phi(R_{i}W_{Q}),\Phi(Y_{j}W_{K})\rangle$. We set $D:=\operatorname{diag}(\beta\cdot A1_{n})\in\mathbb{F}_{p}^{n\times n}$. Then, based on Definition 3.6, we define the i-th kernelized Hopfield layer as $\mathsf{KHop}_{i}(R,Y_{i}):=D^{-1}AY_{i}W_{V}.$
@@
$\mathsf{KHN}(R)=$ $~f_{m}\circ\mathsf{KHop}_{m}(f_{m-1}\circ\mathsf{KHop}_{m-1}(\cdots f_{1}\circ\mathsf{KHop}_{1}(f_{0}(X),Y_{1})\cdots,Y_{m-1}),Y_{m}),$
@@ 4 Complexity of Modern Hopfield Networks
In Section 4.1, we provide some fundamental definitions from circuit complexity theory. This section explores key results regarding the circuit complexity of the computations involved in Modern Hopfield Networks. Then, we begin with an analysis of the Hopfield attention matrix in Section 4.2. In Section 4.3, we delve into the computation of a single Hopfield layer. Section 4.4 extends the discussion to other components beyond the Hopfield layer. In Section 4.5, we examine the modern Hopfield network. Finally, Section 4.6 presents our main result: the circuit complexity bounds for Modern Hopfield Networks. These findings form the basis for our main theorem on the expressive power of Hopfield networks.
4.1 Background on Circuit Complexity
Definition 4.1 (Boolean circuit).
Definition 4.2 (Circuit family and language recognition, Definition 1.10 on page 10 in [45]).
A language $L\subseteq\{0,1\}^{*}$ is a set of binary strings that represent the “yes” instances of a given decision problem, where $\{0,1\}^{*}$ denotes the set of all finite binary strings of any length.